OpenStudy (anonymous):

A teacher gave her class two exams; 60% of the class passed the second exam, but only 48% of the class passed both exams. What percent of those who passed the second exam also passed the first exam?

OpenStudy (anonymous):

80% 12% 30% 48%

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

I got 80% ??

OpenStudy (mathmate):

correct! P(B)=0.60 P(A\(\cap\)B)=0.48 P(A|B)=P(A\(\cap\)B)/P(B)=0.48/0.60=0.80=80%

OpenStudy (anonymous):

Thanks!

OpenStudy (mathmate):

You're welcome! :)